2019 Dec 06 UT: E night

OSA: Josh Williams
Observers: John Hill, Olga Kuhn
SW Support from Igor Sola, and also Petr Kubanek and Dan Cox for the AGw4 problem.

OSCO

canceled as we would have needed extra support for SX/SOUL OSCO

Transform Data collection/verification

on-hold pending a resolution to the AGw4 issue

00:50 Point and collimate near M52.

02:00 Sent L1 script to M52. But we were authorized for bino, so preset failed with co-pointing limit.

Rotator Zeropoints at the start:
LEFT/LUCI RIGHT/LUCI
-187.765 176.85
*John noted, although PCSInstrument.conf says -187.315, the PCS had not been restarted, so PCS thought the rotator zeropoint was -187.765, not -187.315.

02:01 Trying again, after Josh re-authorized for L1 only.
  • L1 images: (sky,obj,mask) = (3,4,5)
  • Results of alignment: dX, dY, dtheta = 6.79, -0.10, -0.42 deg. So the delta theta is -0.42 deg when it was 0 in September and +0.46 for the science masks aligned in October, November and December?
  • Using stars 30-37, dX, dY, dtheta = 6.71, -0.11, -0.40 deg.
  • Mask to detector:
X-shift Y-shift Angle
27.11 -18.1026 0.0981
  • Above was on RM507-5.

02:30 Josh restarted PCS, so the LEFT RZP should be -187.315 now.

02:33 Sending M52 script again.
  • L1 images: (6,7,8)
    • seeing has blown up to 2"
  • Mask to detector:
X-shift Y-shift Angle
25.3703 -17.6464 0.1033
  • Results: initially almost all of the stars have been mis-identified. Taking a while to ID them... Probably the wrong guide star was selected. The images are shifted up by about 100 pixels, but I'm having a hard time finding the stars quickly and we had to retake the acq images after the HBS was restarted, anyway.
  • HBS failed.

02:38 Resending after HBS recovered and pointing checked.
  • L1 images (9,10,11)
  • Mask to detector:
X-shift Y-shift Angle
24.055 -17.55 0.108
  • Results: dX, dY, dtheta = 5.21, -1.283, 0.0455 using all stars (30-38) and with errors 0.101, 0.0998 as.

Transform Data Collection

Doug installed new L2/AGw4 transform. John restarted GCS.

Updating PCSInstrument.conf file in preparation for Transform Data Collection. Since on 20191203, alignments with a 0.45 deg offset resulted still in dtheta=0.01, I gave a 0.46 deg offset this time.

[tcs@tcs1 PCS]$ cp PCSInstrument.conf PCSInstrument.conf.20191205
[tcs@tcs1 PCS]$ vi PCSInstrument.conf
[tcs@tcs1 PCS]$ diff PCSInstrument.conf PCSInstrument.conf.20191203_correct_LUCI1ZP 
203,204c203
< #%LZP changed from -187.315 to -187.775 on 20191206 OK/JMH
< LEFTZEROPOINT      -187.775
---
> LEFTZEROPOINT      -187.765

Rotator Zeropoints are now:
LEFT/LUCI RIGHT/LUCI
-187.775 176.85
IDL> iif_register,/TCS
IDL> iif_preset, Stone_A3, MODE=ACTIVE, GS=0, TS=0, PA=10, /RIGHT
IDL> iif_preset, Stone_A3, MODE=ACTIVE, GS=0, TS=0, PA=10, /LEFT 
IDL> iif_binoc_preset

Configuring LUCIs:
  • N3.75
  • N30_FieldStop
  • Mirror
  • clear+H
  • FlexComp =OFF
Test exposures after Josh stopped guiding & homed guide probes:
  • DIT = 2.51 and NDIT=1, L1: (0012) 52k but L2 (0002) 32k and not in focus on star. Pinholes appear in focus, though.
  • Reinit L2.
  • Repeat L2 (0003) but now LUCI is obviously out of focus. I did an init, but did not reset anything.
  • Toggle filter from Ks to H and repeat (0004). LUCI in focus, star is 39k and a bit blobby. We haven't collimated for a while.

The list I'd chosen, Stone_A3.list, was for MODS and had a few guide stars with numbers exceeding the number of stars in Stone_A3. John and I made a new list, /home/lbto/data/20191206/Stone_A3new.list

03:22 Starting transform_collect, seeing 0
L1 L2 comments
13 5 ok
14 6 background
15 7 gs 11 - L1, ok. but probe failed to move, L2
    gs 12 - L1 did not find star
03:30 GS 11 looks fine. GS 12 is a bit faint (14-16), so maybe that is why. Stone A is very close to the moon.

03:33 Starting over.

03:37 Probe cannot move to requested position (x,y) -0.05, 0.18 mm.

IDL> transform_collect, exp=2.5
  
      IRCR: preset, Stone_A3, MODE=ACTIVE, GS=0, TS=0, PA=-10.0000     (iif_preset)
  
           448 nm rms error in image right_wfscimage000024.fits
Changing sign of Zernikes 5, 8 and 10 for AGw3/AGw4
 tmp_pixels =            1
   => Collimated in 1 iteration
  
      IRCR: preset, Stone_A3, MODE=GUIDE, GS=0, TS=0, PA=-10.0000     (iif_preset)
  
   => guiding stopped
   => GS setting didn't succeed!
   => AGw command status: AGW4: Error moving the motor(s)
   => error: setProbePosition: failed moving probe to position (SFP) (x,y,z): 0, 150, 25.157 mm. Patrol field limits ?
   => problem moving the AGW probe onto the target position
   => AGW4: Error moving the motor(s)
   => Requesting LUCI image
   => Collected //dataB/readout2/DATA/20191206/luci2.20191206.0007.fits
  
      IRCR: offset, 20, 0, ROT=0, COORD=DETXY, MOVE=ABS     (iif_offset)
  
   => Requesting LUCI image
   => Collected //dataB/readout2/DATA/20191206/luci2.20191206.0008.fits
  
      IRCR: preset, Stone_A3, MODE=ACTIVE, GS=0, TS=0, PA=-10.0000     (iif_preset)
  
===> Error in irc_presetTelescope
        err = Resmsg[0]  = PresetTelescope result status: Error
        err = Resmsg[1]  = Stop{FLAO}[459]
        err = Resmsg[2]  = AGw command status: AGW4: Error moving the motor(s)
        err = Resmsg[3]  = isPosAccessible: probe cannot move to requested position (SFP) (x,y): -0.0524156, 0.183271 mm.
        err = Resmsg[4]  = guidestar list successfully changed but failed moving probe to position.

I exited IDL and re-entered and then reissued the "iif_register, RFBG" command. But we did not re-try the transform_collect as we needed to find another list, anyway. Realized that the scale was larger than we thought.

03:53 We are investigating why the AGw4 probe failed to move (the previous error messages).

03:56 Taking a pair of LUCI darks: L1 0020, L2 0009

04:00 Josh restarted both GCS left & right and sending preset again.
  • PA = 0, but error:
  • Fri Dec 6 03:58:28.334 2019 right Preset alert for LUCIFER mode ACQUIRE: isPosAccessible: probe cannot move to requested position (SFP) (x,y): 0.0523898, -0.187636 mm.
  • Removed transform, but still an error. Probe cannot move to requested position (SFP) (x,y): -0.0628907, 0.188863 mm. This is on-axis and should not be a problem.
  • failed moving the motors on "localhost". Transform is absolved, but we don't know what the problem is.
  • restarting OACcontrol, but same error. Josh calling Igor, who is the on-call SW support.
  • Igor, Josh, John troubleshooting. The problem is both with AGW1 and AGw4 and sounds like a computer problem.
  • Calling Petr.

05:00 Switching to LBCs while we continue troubleshooting. No, we went to MODS instead for D M33-1.

D-time MODS observation

completed M33-1 in excellent conditions, both seeing and transparency
observed 1 hr N2403-4 in excellent seeing and cirrus, which caused ~0.5-mag drops for most of the time, but up to 1.5 mag at times during 4th exposure

etaCar analogs OT program file and scripts were placed in ~/Service_Observing/Eng_Dec_2019/etaCar_analogs/

05:37 acqBinoMODS M33-1_blindoff_acq_PA82.txt

MODS2 Offset Command: offsetxy 3.764 7.753 rel
MODS1 Offset Command: offsetxy -0.654 11.622 rel

This left the offset star a bit to the right on MODS1 and pretty good, but slightly too far to the left on MODS2.
dx = -0.123 on SX and dx = 0.1 on DX, to give exposures 0005 and 0005. Need a bit more tweaking.
dx = -0.123 on SX and dx = -0.1 on DX to give exposures 0006 and 0006.

06:00 Petr called. Status now, AGW1 ok, but AGw4 thought to be fixed, but failed on homing.

06:00 Images 0006 on both mods look very well centered. Continuing script to send sync blind offset.

06:01 execBinoMODS M33-1_red_obs.txt to start the 4x900-sec series of red grating exposures.
  • seeing ~0.65" on the guider, down to 0.5.
  • parallactic angle should have remained at 70 deg. I read the wrong axis of the elevation plot, unfortunately, to set it to 82 deg. But we are working in the red where the <~10 deg difference (82 vs 75.5deg, what the par angle is now, 06:38) will be negligible.
  • Spectra are: mods[1|2]r.20191206.0007-0009

06:45 Petr called back saying that he thinks the r-axis is stuck. He can move it with set xy commands, but he cannot move it beyond y=550.

07:00. Clouds coming. Just starting last exposures (4th of 4). We will stop as seeing was very good for the first and we have about 1 mag of clouds at this direction.

07:00 acqBinoMODS N2403-4_PA55_acq.txt
  • acq images are: 10-11 for both mods1r and mods2r.
  • Unfortunately, I sent the direct acq script. Object was very hard to ID.
  • Sending blind offset...created a new one with PA=43 to match par angle.

07:27 acqBinoMODS N2403-4_acq_PA43_blind.txt
  • acq images are: 0012, 0013
    • MODS1 Offset Command: offsetxy -0.628 10.586 rel
  • I sent the command twice for MODS1, but the offsets were different:
    • MODS1 Offset Command: offsetxy -0.641 8.584 rel (because I used -y 9 this time and not -y 11 as before, thinking I was doing the mods2 alignment... but I forgot to change the image names).
  • backed out the last offset, but will retake the field image for MODS1 and use it for alignment.
    • mask is 0014 and field is 0016
      • MODS1 Offset Command: offsetxy -0.090 0.047 rel
    • MODS2: mask is 0014 and field is 0015
      • MODS2 Offset Command: offsetxy 3.925 6.386 rel
    • MODS1 confirmatory image is 0017. Looks good.
    • MODS2 confirmatory image is 0016. The collimator moved? Retaking it. 0017 looks good.
    • Sending blind offset now
08:10 Starting series of science spectra.
  • seeing is 0.8-0.9" on the guider, but clouds are coming and extinction was up to almost 1 mag earlier. It is much less now.
  • Science spectra are: 0018-0021
    • mods2r.20191206.0019-0020 were both corrupted upon copy (as reported in IT). We'll need to recopy it into the archive from mods2data tomorrow (Friday) after noon.
    • On first spectra, I don't see anything very point-like on the slit, though there is a bright-ish object where the target is expected to be.
  • Transparency was pretty good since the sci spectra were started, ~0.25 mag drop at most until just recently (210 sec left of the last exposure). The seeing was excellent. Going to stop after these 4 rather than repeat, as conditions only seem to be deteriorating.
08:18 acqBinoMODS G191-B2B_acq.txt
  • field images are 0022
    • MODS1 Offset Command: offsetxy -0.614 11.913 rel
    • MODS2 Offset Command: offsetxy 4.437 7.710 rel
  • confirmatory images 0023. A bit offset on mods2.
    • MODS2 Offset Command: offsetxy -0.647 -0.254 rel
    • MODS1 Offset Command: offsetxy -0.253 0.008 rel
  • starting spectra - red-only. Some clouds, but here only ~few 0.1s mag extinction, it seems.

Bino LBC superfocs

09:43 reconfiguring for Bino LBCs
  • focus/collimate on field of Eskimo Nebula, at elev 78 deg now and setting.
  • collimated in 3 FPIA iterations after WRS although both had large focus offsets (possibly because temps are on the high side for the winter position, at 4.4C).
  • finished collimation 10:10
  • copointing

10:17 superfoc at PA=0 and Vr

10:26 superfoc at PA=180 and Vr

10:35 superfoc at PA=0 and Vr

10:41 dofpia started again

  -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  2055    -57    104     26   -234    -19     15    y  0.83  0.82                      /newdata/lbcr.20191206.104227.fits  r-SLOAN 
    18    -29     23    -63    205    -24     28    y  0.81  0.13                      /newdata/lbcr.20191206.104355.fits  r-SLOAN 
   624      0      0      0      0   -356      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   829   -172    -75    126     24    -65     14    y  1.09  0.35                      /newdata/lbcb.20191206.104231.fits  V-BESSEL
  -118   -184    -55    -90     49    -30     26    y  1.16  0.12                      /newdata/lbcb.20191206.104359.fits  V-BESSEL
   885      0      0      0      0   -505      0    y  0.00  0.00                     dofpia_sph3_adjustment             

10:45 superfoc PA=0 Vr

10:52 superfoc PA=180 Vr

10:59 superfoc PA=0 Vr

11:05 superfoc PA=90 Vr (clouds are encroaching)

11:13 superfoc PA=270 Vr

11:22 dofpia --- looks like there might have been a seeing bubble at this time, from 1 - 1.5" --- see the DIMM plot.

 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
    99    -92    -11   -172     23     14     11    y  1.32  0.12                      /newdata/lbcb.20191206.112351.fits  V-BESSEL
  -334    -33   -144     95     84    -23      7    y  1.35  0.17                      /newdata/lbcb.20191206.112516.fits  V-BESSEL
   439     16    148    -37   -162    -91      0    y  1.18  0.22                      /newdata/lbcb.20191206.112641.fits  V-BESSEL
  -486   -214   -227    -93    277    -26      0    y  1.26  0.26                      /newdata/lbcb.20191206.112806.fits  V-BESSEL
  -324   -153   -117     19    139     48      0    y  1.32  0.17                      /newdata/lbcb.20191206.112932.fits  V-BESSEL
   985      0      0      0      0   -563      0    y  0.00  0.00                     dofpia_sph3_adjustment       
  -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -611    -80   -127    -80    -64     39     17    y  0.91  0.26                      /newdata/lbcr.20191206.112347.fits  r-SLOAN 
  -660    -14   -207     80     81    106     18    y  0.87  0.29                      /newdata/lbcr.20191206.112512.fits  r-SLOAN 
   726     42     88     -1     64    -92      0    y  0.74  0.31                      /newdata/lbcr.20191206.112637.fits  r-SLOAN 
 -1049    138   -153     11   -198    132     20    y  0.93  0.44                      /newdata/lbcr.20191206.112801.fits  r-SLOAN 
    88    -29    -68      3    -22    -47    -11    y  0.79  0.09                      /newdata/lbcr.20191206.112927.fits  r-SLOAN 
   614      0      0      0      0   -351      0    y  0.00  0.00                     dofpia_sph3_adjustment    

11:30 PA=0 Vr --- first images in this series (extra-focal) are oddly shaped, not donuts. Red looks astigmatic, almost double.

11:38 PA=180 Vr

11:47 PA=0 Vr

dofpia --- DIMM seeing just skyrocketed, from 1.0" at 11:48 UT to 2.2" at 11:56 UT, then back down to 1" at 11:58. FPIA's estimated seeing values reflect the trend.

 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   712    116   -250    -52   -148    153     25    y  1.49  0.33                      /newdata/lbcb.20191206.115424.fits  V-BESSEL
   792    146    145    -42   -204     17     12    y  1.44  0.34                      /newdata/lbcb.20191206.115553.fits  V-BESSEL
   -12    -65   -188     68   -256    -24     -9    y  1.08  0.15                      /newdata/lbcb.20191206.115722.fits  V-BESSEL
    39    279     83    -29     67    -60    -15    y  1.20  0.13                      /newdata/lbcb.20191206.115848.fits  V-BESSEL
   518    -82     37    -28    219   -146    -31    y  1.04  0.26                      /newdata/lbcb.20191206.120013.fits  V-BESSEL
   811      0      0      0      0   -463      0    y  0.00  0.00                     dofpia_sph3_adjustment 
          
 -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1459    -87     -9     37      8     36      1    y  0.83  0.58                      /newdata/lbcr.20191206.115421.fits  r-SLOAN 
   534   -153     96    -29     50     67     23    y  1.24  0.24                      /newdata/lbcr.20191206.115549.fits  r-SLOAN 
   441    -13     59    -99     -3   -163    -20    y  0.83  0.22                      /newdata/lbcr.20191206.115718.fits  r-SLOAN 
   257     61   -178     33   -166   -125    -32    y  0.81  0.18                      /newdata/lbcr.20191206.115844.fits  r-SLOAN 
   -78     33     22     44    253    -51     -2    y  0.68  0.15                      /newdata/lbcr.20191206.120008.fits  r-SLOAN 
   545      0      0      0      0   -311      0    y  0.00  0.00                     dofpia_sph3_adjustment 

12:03 Starting science-like script on Eskimo Nebula using Us, g and r (LBCB) and
  • 12:20 - both blue and red images are elongated. Blue IQ is poor (see 122000)
  • long exposures are very saturated, but check guiding.
  • Seeing has beeen up and down on short timescales.
  • Some images good and some poor. Change of filter? or Z4 corrections?

12:37 dofpia again

 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -2980     60   -314   -171     70      0     23    y  0.88  1.18                      /newdata/lbcb.20191206.123853.fits  V-BESSEL
     0    -38   -267     59     -5    -61    -18    y  1.07  0.12                      /newdata/lbcb.20191206.124024.fits  V-BESSEL
   831      0      0      0      0   -474      0    y  0.00  0.00                     dofpia_sph3_adjustment    
   
  -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   148    -67     10   -347   -233    -80     -5    y  0.56  0.22                      /newdata/lbcr.20191206.123843.fits  r-SLOAN 
  -318    -26    -72    103    -60    -13      3    y  0.66  0.16                      /newdata/lbcr.20191206.124020.fits  r-SLOAN 
   532      0      0      0      0   -304      0    y  0.00  0.00                     dofpia_sph3_adjustment             

12:43 Run OB on same field, but taking a series of short exposures with the Bessell filters. Start out OK, but IQ is deteriorating throughout.

13:01 At 1306, pause and set exptime scaling to 2.0

13:17 Cycle through the filters and take "flats". Lots of clouds. Using the script in ~/Calib_OBs/ClosedDomeTests called allfilt_lowerfloods_PA0 which cycles through all 6 B and 8 R filters and takes a 1-sec exposure in each. Scaled by 3, but stopped and scaled by 10.

MODS Calibrations

mods?r.20191206.00NN.fits

Type of Calibration mods1r mods2r  
Red-Only Slit Flats 0.8" 33-38 33-38  
  Red-Only Slit Flats 1.0" 39-44 39-44  
Slitless Pixel Flats 45- 45-  
Comp lamps: Ne 50 50-51 For MODS2, 51 has no light, 52 is OK. No warm up time has been given for Ne before. Will need to look into this.
Kr+Xe 51 52-53 MODS2 script is taking 2 of each
Ar 52 54-55
Biases 8Kx3K 53-57 56-60

-- OlgaKuhn - 06 Dec 2019
Topic revision: r4 - 07 Dec 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback